### v1

@@ 8 Effects of Superfund remediation on birth outcomes
(b) Spatial weights $w_{1},\ldots,w_{n}$.
(c) $18$ hidden covariates — the nine highest-eigenvalue eigenvectors from both $\bm{S}^{\text{CAR}}$ and $\bm{S}^{\text{GP}}$ .
@@ 9 Practical considerations
We recommend that researchers examine the components of the spatial contrast to understand how information is shared across space and assess whether this aligns with subject-matter knowledge or intuition. To support this, we conclude by describing how borrowing information across space involves navigating a set of tradeoffs.
@@
First, borrowing strength across space reflects a fundamental bias–variance tradeoff. At one extreme, imposing tight balancing constraints on the hidden spatial covariates restricts comparisons to the nearest treated and control units, so information is borrowed only locally. At the other extreme, relaxing these constraints entirely disregards the spatial adjustment performed by the hidden covariates and borrows information globally. The former approach reduces bias from the targeted form of unmeasured spatial confounding at the cost of greater variance, whereas the latter approach lowers variance but is more susceptible to unmeasured spatial confounding bias. This result mirrors the analytical result in 11.2.1 and aligns with the intuition behind donor pool selection in synthetic control analyses (Abadie 2021). We therefore recommend that investigators (i) carefully examine the distribution of the weights in space and (ii) assess the sensitivity of point estimates and standard errors across a plausible range of balancing thresholds for the hidden covariates.
@@
Second, borrowing strength across space introduces a tension between adjusting for unmeasured spatial confounding bias and avoiding bias from spatial interference. When interference is present, using the most proximal control units as comparisons may be inappropriate, as these units could be affected by the treatment received by nearby treated units. While we do not address spatial interference in the current work, previous studies have proposed selecting control units beyond a spatial buffer, so they are not exposed to the treatment but still have similar values of the unmeasured spatial confounder (Butts 2021). We recommend that, in settings where interference is plausible, researchers consider such buffered comparisons to reduce potential contamination.
@@
A third tradeoff introduced by borrowing strength across space relates to the positivity assumption and the target population. When borrowing strength only locally, some treated units may be too distant from any control units to permit meaningful comparisons. In such cases, we argue that these treated units should be excluded from the target population, as there is no justification for assuming the existence of control units with similar values of the unmeasured confounder (a violation of positivity). Thus, imposing strict balancing constraints on the hidden covariates may require us to redefine the target population. In light of this, we recommend that researchers carefully consider the choice of target population before proceeding with any analysis. All of these comments are in the spirit of matching and weighting methods for observational studies (Zubizarreta et al. 2023). Embedding spatial-contrast considerations into routine practice will strengthen the credibility and interpretability of spatial causal analyses.
@@ 10 Discussion
In this paper, we unify three types of spatial regression through a weighting framework for causal inference. Our exact, finite-sample analysis reveals that spatial regression models adjust for unmeasured spatial confounding by approximately balancing a hidden set of spatially autocorrelated covariates determined by the spatial error structure. Furthermore, we show that the confounding bias due to an unmeasured confounder $U$ is small whenever $U$ has high Moran’s I. Our mathematical characterization is accompanied by three visual diagnostics to better understand this result. Finally, we propose a more general spatial weighting estimator that accommodates nonlinearities and treatment effect heterogeneity in the outcome model and simultaneously adjusts for unmeasured confounders with different forms of spatial autocorrelation.
@@
From a regression modeling perspective, our main result makes explicit how the assumed error structure in spatial regression adjusts for unmeasured confounders. Specifically, we show that incorporating a spatial error term is equivalent to augmenting the linear model with additional auxiliary covariates (the eigenvectors of the covariance matrix) whose coefficients are penalized, thereby inducing an approximate mean balance on these covariates. In this sense, “modeling the error” and “softly balancing eigenvectors of the covariance matrix” are two sides of the same coin. More broadly, under a linear model, the error covariance structure can be re-expressed as additional covariate structure with appropriate regularization, making these implicit adjustments explicit and more interpretable.
@@
We highlight two avenues for future work. First, although we provide an exact, finite-sample analysis that informs large-sample considerations, the development of a comprehensive asymptotic theory remains open. Gilbert et al. (2025) established consistency for Gaussian process regression, and Datta and Stein (2025) derived similar results for settings where the exposure, outcome, and unmeasured confounder are generated by Gaussian random fields, both under infill asymptotics. However, extending this theory to random effects or conditional autoregressive models presents additional challenges that warrant further investigation. As $n$ grows, new units may join existing clusters and form new adjacencies, so pre-existing entries of the covariance matrix change with sample size. Accounting for this evolving geometry is essential to any asymptotic treatment of these models.

### v2

@@ 7 Effects of Superfund remediation on birth outcomes
(b) Thirty latent covariates—the ten highest-eigenvalue eigenvectors from $\bm{S}^{\text{RE}}$, $\bm{S}^{\text{CAR}}$, and $\bm{S}^{\text{GP}}$.
(c) Spatial weights $w_{1},\ldots,w_{n}$.
@@ 8 Practical considerations
In general, we recommend that researchers examine the weights of the treated and control components of the spatial estimator to understand how information is used across space and to assess whether this pattern aligns with subject-matter knowledge or intuition. To support this, we conclude by describing how borrowing information across space involves navigating a set of interrelated tradeoffs.
@@
First, borrowing strength across space reflects a fundamental bias–variance tradeoff. At one extreme, imposing tight balancing constraints on the latent covariates restricts comparisons to the nearest treated and control units, so information is borrowed only locally. At the other extreme, relaxing these constraints effectively disregards the spatial adjustment performed by the latent covariates and borrows information globally. The former approach reduces bias from the targeted form of unmeasured spatial confounding at the cost of increased variance, whereas the latter approach lowers variance but is more susceptible to unmeasured spatial confounding bias. This result mirrors the analytical result in Supplementary Section B and aligns with the intuition behind donor pool selection in synthetic control analyses (Abadie 2021). We therefore recommend that investigators (i) carefully examine the distribution of the weights in space and (ii) assess the sensitivity of point estimates and standard errors across a plausible range of balancing thresholds for the latent covariates.
@@
Second, borrowing strength across space introduces a tension between adjusting for unmeasured spatial confounding bias and avoiding bias from spatial interference. When interference is present, using the most proximal control units as comparisons may be inappropriate, as these units could themselves be affected by the treatment received by nearby treated units. Previous studies have proposed selecting control units immediately outside a spatial buffer, so they are not treated to the treatment but still have similar values of the unmeasured spatial confounder (Butts 2021). In settings where interference is plausible, we recommend considering such buffered comparisons to reduce potential contamination.
@@
A third tradeoff introduced by borrowing strength across space relates to the positivity assumption and the target population. When borrowing strength only locally, some treated units may be too distant from any control units to permit meaningful comparisons. In such cases, we argue that these treated units should be excluded from the target population, as there is no justification for assuming the existence of control units with similar values of the unmeasured confounder (a violation of positivity). Thus, imposing strict balancing constraints on the latent covariates may require redefining the target population. In light of this, we recommend that researchers carefully consider the choice of target population before proceeding with any analysis. All of these considerations are in the spirit of matching and weighting methods for observational studies (Zubizarreta et al. 2023). Embedding spatial-contrast considerations into routine practice can strengthen both the credibility and interpretability of spatial causal analyses.
@@ 9 Discussion
This paper unifies three classes of spatial regression models within a common weighting framework for causal inference. Our exact, finite-sample analysis reveals that spatial regression adjusts for unmeasured spatial confounding by approximately balancing a latent set of spatially autocorrelated covariates determined by the assumed spatial error structure. Furthermore, we show that the confounding bias induced by an unmeasured confounder $\bm{U}$ is small whenever $\bm{U}$ exhibits strong spatial autocorrelation as measured by Moran’s I. We propose visual diagnostics to better understand this result, as well as a more general spatial weighting estimator that simultaneously adjusts for unmeasured confounders with different forms of spatial autocorrelation, which are typically considered separately (i.e., one at a time) in spatial regression. In an observational study of Superfund site remediation and Medicaid-covered birth outcomes, we find that remediation between 1991–2015 potentially decreased the rate of small vulnerable newborns in nearby areas during 2016–2018 (ATT estimate: 0.60-0.60%).
@@
Our main result demonstrates that incorporating a spatial error term is equivalent to augmenting the linear model with penalized eigenvectors of the covariance matrix. This formulation implicitly induces approximate mean balance on these latent covariates. In this sense, “modeling the error” and “softly balancing eigenvectors of the covariance matrix” are two sides of the same coin. While we develop these results in the context of spatial regression, they broadly apply to linear regression models with generalized error structures. In other words, under a linear model, the error covariance structure can be re-expressed as additional covariate structure with appropriate regularization, making these implicit adjustments explicit and more interpretable.
@@
We highlight two avenues for future work. First, although we provide a finite-sample analysis that informs large-sample considerations, the development of a comprehensive asymptotic theory remains open. Gilbert et al. (2025) established consistency for Gaussian process regression, and Datta and Stein (2025) derived consistency results for settings where the treatment, outcome, and unmeasured confounder are generated by Gaussian random fields, both under infill asymptotics. However, extending this theory to random effects or conditional autoregressive models presents additional challenges that warrant further investigation. As the sample size grows, new units may join existing clusters and form new adjacencies, so pre-existing entries of the covariance matrix change with sample size. Accounting for this evolving geometry is essential to any asymptotic treatment of these models.